A sound wave radiates from a source in all directions. If thepower of the sound source is 200 watts, then the intensity level ofthe sound wave 100m from the source is, in dB
Answer #1
We power   P = Intensity *Area
Area =4πR2
R= 100 m
P = 200W
Then Intensity I= P/Area
I = 200W/4π(100m)2 = 1.592*10-3W/m2
We knowthat intensity level β = (10dB) log(I/Io)
Here Io= 10-12W/m2
Then   β =  (10dB) log(I/Io)
β= (10dB)log( 1.592*10-3W/m2/10-12W/m2) =92dB